OpenStudy (babyslapmafro):

Please help me find the inverse Laplace transform of -2/(s+2)^2

OpenStudy (babyslapmafro):

To clarify, I'm trying to find the inverse Laplace transform of... \[\frac{ -2 }{ (s+2)^{2} }\]

OpenStudy (babyslapmafro):

My answer is... \[-2e ^{-4t}\] but it is incorrect.

OpenStudy (loser66):

On the table of laplace transform, formula 23 tells you that inverse Laplace transform of \(\dfrac{n!}{(s-a)^{n+1}}\) is \(t^ne^{at}\) Your function is \(\dfrac{-2}{(s+2)^2}\) so, your n+1 = 2 or n =1, and a =(-2) It becomes: \(\dfrac{1!}{(s-(-2))^2}\) Hence Inverse of it will be \(t^1 e^{-2t}=te^{-2t}\)

OpenStudy (loser66):

Sorry, I forgot -2 on numerator, so, the answer should be \(-2te^{-2t}\)

OpenStudy (babyslapmafro):

My table has 7 transforms to choose from, I suppose I'll need to find another table.

OpenStudy (babyslapmafro):

Why can't I take the transform for 1/(s+a) and square it to get the answer? This is what I did to get my original answer.

OpenStudy (babyslapmafro):

@Loser66

OpenStudy (loser66):

Because you don't have that form.

OpenStudy (loser66):

Go through the definition of Laplace transform to see why you can't use it.

OpenStudy (babyslapmafro):

\[\frac{ 1 }{ s+a } \times \frac{ 1 }{ s+a }\]

OpenStudy (loser66):

You don't have it

OpenStudy (babyslapmafro):

How do I "not have it"?

OpenStudy (loser66):

What do you want me to explain? from formula or from definition or properties of Laplace ?

OpenStudy (babyslapmafro):

Idk, forget it

OpenStudy (loser66):

ok, I go from definition of Laplace \[\mathcal{L}(1)=\int_0^{\infty}e^{-st}dt=\dfrac{1}{s}\] Now, \[\mathcal{L}(f(t))=\int_0^\infty e^{-st}f(t)dt\] So if \[\mathcal{L}(te^{-2t})=\int_0^\infty e^{-st}te^{-2t}dt=\int_0^\infty te^{-t(s+2)}dt\] let u =t, du =dt, \(dv=e^{-t(s+2)}dt\rightarrow v=\dfrac{-e^{-t(s+2)}}{s+2}\) so the integral is \[t\dfrac{-e^{-t(s+2)}}{s+2}(from~0 ~to~\infty)+\int_0^{\infty} \frac {e^{-t(s+2)}}{s+2}dt\]

OpenStudy (loser66):

the first guy gives you 0, the second guys give you \(\dfrac{-1}{(s+2)^2}\)

OpenStudy (loser66):

If you apply definition for your answer, the Laplace transform is not that Thanks for asking. I have a good chance to review.